OpenStudy (anonymous):

Real analysis... Prove that the limit (as n approaches infinity) of (1/n)sin(n)=0

OpenStudy (anonymous):

So far I have \[\left| \frac{ \sin(n) }{n } \right|<\epsilon \] Where do I go from here? Do I split it up into 1/n and sin(n)?

OpenStudy (anonymous):

what do you know about the sin function (for example, is it bounded by something)? we can say that \abs(sin(n)/n)=\abs(sin(n))/\abs(n)